OpenStudy (anonymous):

Factor completely 4x3 + 8x2 - 25x – 50. (4x2 - 25)(x + 2) (2x - 2)(2x + 2)(x + 5) (2x - 5)(2x + 5)(x + 2) (2x - 5)(2x + 5)(x - 2)

OpenStudy (anonymous):

Given the polynomial 2x3 + 18x2 - 18x - 162, what is the value of the coefficient 'k' in the factored form? 2x3 + 18x2 - 18x - 162 = 2(x + k)(x - k)(x + 9) k= ____________ Numerical Answers Expected! Answer for Blank 1:

OpenStudy (anonymous):

i just need a little help if u want u can just give the answers up to u

OpenStudy (anonymous):

Look at the First one. And Split the Equation in 2. So Factor \[4x^3+8x^2\] And Factor \[-25x-50\]

OpenStudy (anonymous):

Then tell me what you get c:

OpenStudy (anonymous):

4x\[4x^3+8x^2= 4x^2(x+2)\]

OpenStudy (anonymous):

Yess, And do the Samme for the other one. \[4x^2(x+2)-(25x-50)\]

OpenStudy (anonymous):

the 2nd one is\[-25(x-2)\]

OpenStudy (anonymous):

Try again.

OpenStudy (anonymous):

Remember, two negatives make a positive, and it isnt positivee.